The cost for a rental car is $49.95 per day plus a flat fee of $31.95 for insurance. The equation, C=49.95x+31.95 represents the total cost, C (in dollars), to rent the car for x days. Identify the slope and interpret the meaning of the slope in the context of this equation. Identify the C-intercept and interpret the meaning of the C-intercept in the context of this equation. How much would it cost to rent the car for 1 week?
Now that you've used that example of an equation in slope-intercept form, please give an example of another calculation similar to this.

Answer by Theo(13342)

the equation is c = 49.95 * x + 31.95

this is in the form of the straight line equation of y = mx + b
m is the slope.
b is the y-intercept.

in the equation of c = 49.95 * x + 31,95, ....

the slope is 49.95 and the c-intercept is 31.95.

the c-intercept is the cost when x = 0.
this cost is paid regrdless of the number of days the car is rented.

the slope is the incremental cost per day.
for 1 day, it's 49.95
for 2 days, it's 49.95 * 2 = 99.90
for 3 days, it's 49.95 * 3 = 149.85
etc.....

the total cost for 1, 2, 3 days would be that plus 31.95.

for 1 day, it's 45.99 + 31.95 = 77.94
for 2 days, it's 99.90 + 31.95 = 123.93
for 3 days, it's 149.85 + 31.95 = 169.92
etc.....

replace c with y and you can graph the equation as shown below.



if you rent the car for 1 week, that's 7 days.

the total cost will be 45.99 * 7 + 31.95 = 353.88.

another calculation similar to this might be:

rent a boat for 99.99 a day plus insurance of 100.00.

let c = the total cost and x be the total number of days.
the equation is c = 99.99 * x + 100.
99.99 is the slope which is the incremental cost per day.
100 is the one time up front cost of 100 for the insurance.